# LGR2: Language Guided Reward Relabeling for Accelerating Hierarchical Reinforcement Learning

### v2

@@ 4 Methodology
In this section, we introduce LGR2: Language Guided Reward Relabeling for Accelerating Hierarchical Reinforcement Learning, a novel HRL approach that leverages natural language instructions to solve complex sparse reward robotic tasks. The main idea behind our approach is to translate language instructions to parameters for the higher-level reward function $r_{\phi}$, which is subsequently used to relabel higher-level replay buffer transitions. Since $r_{\phi}$ is independent of lower primitive behaviour, this mitigates reward non-stationarity (Section 3.1.1). However, in sparse reward environments, $r_{\phi}$ may fail to generate a meaningful reward signal. To resolve this issue, we additionally incorporate hindsight relabeling to reduce reward sparsity and enhance the informativeness of the language-generated reward function. Thus, LGR2 is an efficient language-guided HRL framework that translates language instructions to robot actions, while mitigating non-stationarity in HRL.
The remainder of this section is organized as follows. First, we present our language-guided reward relabeling procedure to mitigate non-stationarity in HRL. We then explain our goal-conditioned hindsight relabeling approach for reducing reward sparsity. Finally, we conclude by introducing the practical implementation and an overview of our algorithm LGR2.
4.1 Language Guided Reward Relabeling
In our goal-conditioned HRL framework, the higher-level policy $\pi^{H}$ executes for $n$ timesteps, and provides subgoals $g_{t}$ to the lower-level policy $\pi^{L}$, which executes for $k$ timesteps between each consecutive pair of subgoal predictions. The higher-level policy $\pi^{H}$ stores samples of the form $(s_{t},g^{*},g_{t},r^{H}_{t}=\sum_{i=t}^{t+k-1}r^{L}_{i},s_{t+k})$ in its replay buffer, where higher level reward $r^{H}_{t}$ is the sum of environment rewards encountered by the lower level policy for $k$ timesteps. Due to the non-stationary behaviour of the lower primitive, the higher level reward $r^{H}_{t}$ might become outdated. The key underlying idea of our approach is to relabel the reward function $r^{H}_{t}$ using the reward parameters $r_{\phi}$ generated using language instructions. Concretely, first the language instructions are translated to parameters of the higher level reward function $r_{\phi}$ using L2R [19] and subsequently, the higher level replay buffer transitions are relabeled using $r_{\phi}$, to yield the following form: $(s_{t},g^{*},g_{t},r_{\phi},s_{t+k})$. Note that it is straightforward to extract the final goal parameters from $r_{\phi}$ using L2R motion descriptor, since the parameters specify the final goal configuration in goal-conditioned RL. We formulate the motion descriptor and reward coder prompts in the reward translator module in L2R according to our robotic navigation and manipulation tasks. The reward translator prompts are provided in detail in the Appendix A.3. Since $r_{\phi}$ remains unaffected by lower primitive behaviour, this helps alleviate the reward non-stationarity discussed in Section 3.1.1.
Although, combining the advancements in L2R and HRL does provide the benefits of language-guided robotic control, temporal abstraction, and non-stationarity mitigation, however, in our HRL setup, we found in practice that the higher level reward function $r_{\phi}$ generated by L2R is too sparse for the lower primitive to make any substantial progress. To this end, we next propose an approach for densifying higher level rewards generated from L2R from language instructions.
@@ 4.1 Language Guided Reward Relabeling
11:         $\sigma\sim\mathnormal{R^{H}}$ where $\sigma=\{(s_{t},g^{*},g_{t},\sum_{i=t}^{t+k-1}r_{i},s_{t+k-1})\}_{t=1}^{n-1}$
@@ 4.2 Goal-conditioned Hindsight Relabeling
Although the language-guided higher-level reward $r_{\phi}$ presented in Section 4.1 is intuitively appealing for relabeling higher-level replay buffer transitions, its sparsity significantly undermines its ability to produce a meaningful reward signal. To address this issue, we employ hindsight relabeling (HER) [21] for densifying the language-guided rewards. Concretely, we employ HER to generate additional replay buffer transitions by relabeling the goal and rewards.
@@ 4.3 LGR2 Implementation
We explain our approach in detail in Algorithm 1. The higher and lower-level policies undergo training using SAC [72]. The rewards in higher level transitions are relabeled using the language-guided reward $r_{\phi}$. Rather than relabeling all transitions in the replay buffer, we selectively relabel transitions from the higher-level replay buffer as they are sampled during training.
@@ 5 Experiments
Figure 2: Success rate comparison This figure compares the success rate performances on four sparse maze navigation and robotic manipulation environments. The solid line and shaded regions represent the mean and standard deviation, across $5$ seeds. We compare our approach LGR2 against multiple baselines. PIPER shows impressive performance and significantly outperforms the baselines.
@@
In this section, we perform experimental analysis to answer the following questions: (1) How effectively does LGR2 perform in sparse maze navigation and robotic manipulation tasks? (2) Can LGR2 address the recurring issue of non-stationarity in HRL? (3) Does LGR2 outperform single level L2R approach? (4) Does LGR2 improve sample efficiency and training stability? (5) What is the impact of using HER on the performance in LGR2?
@@
Implementation details We evaluate PIPER on four robotic navigation and manipulation tasks with continuous state and action spaces: $(i)$ maze navigation, $(ii)$ pick and place [21], $(iii)$ bin, and $(iv)$ franka kitchen [73]. We provide a video depicting qualitative results and the implementation code in the supplementary. We also provide extensive details about the environments and implementation in the Appendix A.1. For our RL algorithm, we utilize the off-policy SAC [72] with Adam [74] optimizer. To ensure fair comparisons, we maintain consistent training conditions across all baselines, and we empirically tune the hyper-parameters for our method and all other baselines. Additionally, since the pick and place, push and kitchen task environments are difficult, in order to speedup training, we assume access to a single human demonstration, and use an additional imitation learning objective at the lower level. We do not assume access to any demonstration in the maze navigation task. We keep this assumption consistent among all baselines to ascertain fair comparisons.
@@
Evaluation and Results To analyze our design choices, we consider several baselines: L2R [19], HAC (Hierarchical Actor Critic [69]), HIER (vanilla hierarchical SAC), and DAC (Discriminator Actor Critic [75]). In the original L2R approach [19], the language instructions are translated to the parameters of reward function via a reward translator module, and the reward function is optimized using a Mujoco MPC controller [20]. In order to ascertain fair comparisons, we implement the original reward translator module in our L2R implementation, but we replace the MPC controller by an RL agent implemented using SAC [72]. Hence, the reward translator module generates reward function parameters, which are optimized using RL. We compare our approach with this baseline to demonstrate the importance of our hierarchical formulation. As seen in Figure 2, our approach significantly outperforms this baseline, which shows that our hierarchical formulation is crucial for solving complex robotic control tasks that require long term planning. We also compare our approach with HAC baseline, which deals with the non-stationarity issue by assuming the lower-level primitive to be optimal. Concretely, HAC augments the higher-level replay buffer with additional transitions by relabeling the subgoals in the replay buffer while assuming an optimal lower primitive. As seen in Figure 2, our approach surpasses HAC, indicating that our language-guided reward relabeling approach is able to better mitigate non-stationarity issue in HRL. Further, we consider HIER baseline, which is a vanilla hierarchical SAC implementation, where the higher level reward is the sum over k-step environment rewards. Since LGR2 significantly outperforms this baseline, this demonstrates that language-guided reward generation is indeed able to mitigate non-stationarity in HRL. Finally, we also consider DAC baseline with access to a single demonstration. DAC failed to show any significant progress, as seen in Figure 2, which shows that LGR2 is even able to outperform single-level baselines which have access to privileged information like expert demonstrations.
@@
Figure 4: Real world experiments in pick and place and bin environments. Row 1 depicts initial and Row 2 depicts goal configuration.
Ablation Analysis Here, we analyse the importance of using hindsight experience replay (HER) for densifying rewards, when dealing with sparse reward tasks. We compare LGR2 with LGR2-No-HER baseline, which is LGR2 baseline without HER. As seen in Figure 4, HER is crucial for good performance, since language-guided rewards are too sparse to generate any meaningful reward signal.
Real World Experiments We perform real world experiments in real pick and place and bin and environments (Fig 4). We use Realsense D435 depth camera to extract the robotic arm, block and bin positions. We train the hierarchical policies in simulation and deploy on real robotic arm. Additionally, we found that assigning small hard-coded values to the linear and angular velocities shows good performance, when deploying on the real robot. LGR2 achieves an accuracy of $0.6$ and $0.5$ on pick and place and bin environments respectively. We also deployed the next best performing baseline L2R on the two tasks, but it was unable to achieve the goal in any of the tasks.
@@ 6 Conclusion
Limitations and future work Although we demonstrate that LGR2 is able to solve complex robotic control tasks through natural language instructions, it faces certain limitations. Firstly, designing prompts for the reward translator module requires additional manual effort. Alternatively, if we neglect designing the prompts, the LLMs might hallucinate the parameters for reward function. An interesting direction for future work is to automatically generate the prompts, or even disregard the prompts altogether. Further, LGR2 first employs LLMs to relabel higher level rewards, and then subsequently optimizes the rewards using RL. This two-step approach adds an additional layer of complexity. A possible solution is to combine the higher-level policy and reward translator module, where the higher-level policy uses natural language instructions to generate subgoals. However, it is challenging to design subgoals that are not easily expressible using natural language. An interesting research direction is to design efficient intermediate subgoal representations for the lower primitive, which we plan to analyse in the future.
Appendix A Appendix
A.1 Implementation details
We conduct our experiments on two systems, each equipped with Intel Core i7 processors, 48GB of RAM, and Nvidia GeForce GTX 1080 GPUs. We also report the timesteps required for running the experiments. In our experiments, the actor and critic networks are designed as three-layer, fully connected neural networks with $512$ neurons per layer. In the maze navigation task, a $7$-degree-of-freedom (7-DoF) robotic arm navigates a four-room maze, with its closed gripper (fixed at table height) moving through the maze to reach the goal position. For the pick and place task, the $7$-DoF robotic arm’s gripper locates a square block, picks it up, and delivers it to the goal position. In bin environment, the gripper has to pick up the block and place it in a specific bin. In the kitchen task, a $9$-DoF Franka robot executes a pre-defined complex task, specifically opening the microwave door, to achieve the final goal. To ensure fair comparisons, we maintain consistency across all baselines by keeping parameters such as neural network layer width, number of layers, choice of optimizer, and SAC implementation parameters the same wherever possible.

### v3

@@ 4 Methodology
In this section, we introduce LGR2: Language Guided Reward Relabeling for Accelerating Hierarchical Reinforcement Learning, a novel HRL approach that leverages natural language instructions to mitigate non-stationarity and effectively solve complex sparse reward robotic tasks. The key underlying idea behind our approach is to translate the natural language instructions to reward parameters $r_{\phi}$ for the higher-level policy, which is subsequently used to relabel higher-level replay buffer transitions. Since $r_{\phi}$ is independent of lower primitive behaviour, this mitigates reward non-stationarity in HRL. Further, in sparse reward environments, $r_{\phi}$ may fail to generate a meaningful reward signal. To resolve this, we additionally incorporate hindsight relabeling (Andrychowicz et al., 2017) to reduce the reward sparsity of the language-generated reward function.
The remainder of this section is organized as follows: $(i)$ we first present our language-guided reward relabeling procedure to mitigate non-stationarity in HRL, $(i)$ we then explain our goal-conditioned hindsight relabeling approach for reducing reward sparsity, and $(iii)$ we finally conclude by providing a detailed overview of our LGR2 algorithm and its practical implementation.
@@
11:         $\sigma\sim\mathnormal{R^{H}}$ where $\sigma=\{(s_{t},g^{*},g_{t},r^{H}_{t},s_{t+k-1})\}_{t=1}^{n-1}$
@@
4.1 Language to Rewards
In our goal-conditioned HRL framework, the higher-level policy $\pi^{H}$ stores samples of the form $(s_{t},g^{*},g_{t},r^{H}_{t},s_{t+k})$ in its replay buffer, where higher level reward $r^{H}_{t}$ denotes the environment rewards. Due to changing lower primitive, the higher level reward $r^{H}_{t}$ becomes non-stationary. We now define our approach for translating natural-language user instructions into higher-level policy reward functions. Although we can use Large Language Models (LLMs) to directly generate reward functions for simple tasks, it is hard to generate reward functions since we deal with sparse complex long horizon tasks. Hence, inspired from Yu et al. (2023), we define a two-stage approach by first leveraging a motion descriptor module that interprets and expands user prompts into a structured, natural-language description of the target movement and then by subsequently leveraging a reward coder module which translates this detailed description into reward functions.
Motion Descriptor Module. Although the LLMs struggle to directly generate complex reward functions, they excel at describing motion in natural language. Based on this insight, we design a motion descriptor template that captures common robot movements, allowing the LLM to generate structured and predictable task-specific motion description output. Since the motion is described in natural language, the LLM can generate results without requiring specific example prompts.
Reward Coder Module. Subsequently, the reward coder module converts the motion description into a reward function by leveraging its coding knowledge to produce precise, well-structured code. The prompt for the Reward Coder includes three key components: a description of the available reward APIs, an example of the expected output, and specific constraints the LLM must follow. This module determines reward parameters based on its understanding of the motion description, ensuring accurate and context-aware reward generation.
By separating motion description from code generation, we exploit the advantage of LLMs in both natural-language reasoning and coding, making it easier to create domain-specific, customizable reward functions for robotics tasks. Finally, we extract the final goal parameters from $r_{\phi}$ which specify the final goal configuration in goal-conditioned RL. Please refer to the supplementary for both the motion descriptor and reward coder module prompts for all environments.
Language Guided Reward Relabeling The reward parameters $r_{\phi}$ generated using language instructions are subsequently used to relabel the reward function $r^{H}_{t}$. Concretely, the higher level replay buffer transitions are relabeled using $r_{\phi}$, to yield the following form: $(s_{t},g^{*},g_{t},r_{\phi},s_{t+k})$, as shown in Figure 1. Since the reward no longer depends on the changing lower primitive behavior, it mitigates non-stationarity in HRL and stabilizes HRL training. Finally, the higher-level and lower-level policies are trained using reinforcement learning. Thus, LGR2 harnesses the combined benefits of language-guided robotic control, temporal abstraction, and improved exploration while stabilizing training via HRL non-stationarity mitigation.
However, we found that in practice, the higher level reward function $r_{\phi}$ is too sparse for the lower primitive to make any substantial progress. To this end, we next propose an approach for densifying higher-level rewards using hindsight relabeling.
@@ 4.2 Goal-conditioned Hindsight Relabeling
Although the language-guided higher-level reward $r_{\phi}$ presented in Section 4.1 is intuitively appealing for relabeling higher-level replay buffer transitions, its sparsity significantly undermines its ability to produce a meaningful reward signal. To address this issue, we employ hindsight relabeling (HER) (Andrychowicz et al., 2017) for densifying the language-guided rewards. Concretely, we employ HER to generate additional replay buffer transitions by relabeling the goal and rewards.
@@ 4.3 LGR2 Implementation
We explain our approach in detail in Algorithm 1. The hierarchical policies undergo training using SAC (Haarnoja et al., 2018). The rewards in higher level transitions are relabeled using the language-guided reward $r_{\phi}$. Rather than relabeling all transitions in the replay buffer, we selectively relabel transitions from the higher-level replay buffer as they are sampled during training.
@@ 5 Experiments
Figure 2: Success rate comparison This figure compares the success rate performances on four sparse maze navigation and robotic manipulation environments. The solid line and shaded regions represent the mean and standard deviation, across $5$ seeds. We compare our approach LGR2 against multiple baselines. LGR2 shows impressive performance and significantly outperforms the baselines.
@@
In this section, we perform experimental analysis to answer the following questions: (1) How well does LGR2 perform in sparse maze navigation and robotic manipulation tasks, compared to prior baselines? (2) Can LGR2 mitigate non-stationarity in HRL? (3) Does LGR2 outperform single level L2R approach? (4) Does LGR2 improve sample efficiency and training stability? (5) What is the empirical benefit of using HER?
Implementation details We evaluate LGR2 on four continuous robotic navigation and manipulation tasks: $(i)$ maze navigation, $(ii)$ pick and place (Andrychowicz et al., 2017), $(iii)$ bin, and $(iv)$ franka kitchen (Gupta et al., 2019). These are formulated as sparse reward scenarios, where the agent is only rewarded when it comes within a $\delta$ distance of the goal. Due to this, these environments are hard where the agent must extensively explore the environment before coming across any rewards. As an example: in franka kitchen task, the agent only receives a sparse reward after achieving the final goal (e.g. successfully opening the microwave and then turning on the gas knob). Additionally, we consider goal-conditioned RL setting, where the final goals are randomly generated, increasing the task complexity. These nuances prohibit the prior baselines from performing well in these tasks, which makes these test beds ideal scenarios for evaluation. In the maze task, we use intermediate waypoints as subgoals, whereas for complex tasks (e.g. franka kitchen), we consider full intermediate states as waypoints or subgoals. The implementation code for LGR2 and a detailed video depicting qualitative results are provided in the supplementary.
@@
We provide additional implementation and environment details (Supplementary Sections B and C) and prompts and codes used in experiments (Supplementary Sections D and E). The actor and critic networks are designed as three-layer, fully connected neural networks with $512$ neurons per layer. We use SAC (Haarnoja et al., 2018) with Adam (Kingma & Ba, 2014) optimizer. To ensure fair comparisons, we maintain consistency across all baselines by keeping the hyper-parameters consistent across baselines. Additionally, we extensively tune the hyper-parameters for our method and all other baselines. Since the pick and place, push, and kitchen task environments are hard, we assume access to a single human demonstration and use an additional imitation learning objective at the lower level to speed up training. We do not assume access to any demonstration in the maze navigation task. We keep this assumption consistent among all baselines to ascertain fair comparisons.
Evaluation and Results We compare LGR2 against multiple hierarchical and non-hierarchical baselines. Please refer to Figure 2 for success rate comparison plots and subsequent discussion. The solid line and shaded regions represent the mean and standard deviation, averaged over 5 seeds.
Comparison against L2R baseline.
@@
In the original L2R approach (Yu et al., 2023), the language instructions are translated to the parameters of reward function via a reward translator module, and the reward function is optimized using a Mujoco MPC controller (Howell et al., 2022). In order to ascertain fair comparisons, we implement the original reward translator module in our L2R implementation, but we replace the MPC controller by an RL agent implemented using SAC (Haarnoja et al., 2018). Hence, the reward translator module generates reward function parameters, which are optimized using RL. We compare our approach with this baseline to demonstrate the importance of our hierarchical formulation. Notably, since we deal with complex sparse reward scenarios, the baseline fails to show significant progress. As seen in Figure 2, our approach significantly outperforms this baseline, which shows that our hierarchical formulation is crucial for solving complex robotic control tasks that require long term planning.
Comparison against hierarchical baselines.
SAGA. We compare LGR2 with SAGA (Wang et al., 2023), a hierarchical approach that uses state-conditioned discriminator network training to mitigate non-stationarity. SAGA ensures that the high-level policy generates subgoals aligned with the current state of the low-level policy. While SAGA performs well in maze environment, it struggles with more challenging tasks where LGR2 significantly outperforms it. This suggests that SAGA is more susceptible to non-stationarity in complex, long-horizon tasks, whereas LGR2 is more effective at addressing this issue.
RAPS. The RAPS baseline (Dalal et al., 2021) uses a library of robot action primitives parameterized with arguments that are learned by RL. This baseline uses pre-defined control primitives directly controlled by the higher level policy. RAPS outperforms LGR2 in the maze environment, which shows that approaches that employ pre-defined primitives might be good choices for solving relatively simpler environments. However, we find that it fails to show progress in the harder sparsely rewarded environments (pick and place, bin, kitchen). This demonstrates that there is need for developing end-to-end methods like LGR2 that employ the inherent hierarchical structure for solving complex long-horizon and multi-step tasks.
HAC. We compare our approach with the HAC (Andrychowicz et al., 2017) baseline that deals with the non-stationarity issue by assuming the lower-level primitive to be optimal. Concretely, HAC augments the higher-level replay buffer with additional transitions by relabeling the subgoals in the replay buffer while assuming an optimal lower primitive. As seen in Figure 2, our approach surpasses HAC, indicating that our language-guided reward relabeling approach is able to better mitigate non-stationarity issue in HRL.
HIER. Further, we consider HIER baseline, a vanilla hierarchical SAC implementation, where the higher level reward is the sum over k-step environment rewards. Since LGR2 significantly outperforms this baseline, this demonstrates that language-guided reward generation is indeed able to mitigate non-stationarity in HRL.
Comparison against non-hierarchical DAC baseline.
Finally, we also consider DAC baseline with access to a single demonstration. DAC fails to show any significant progress, as seen in Figure 2, which shows that LGR2 outperforms single-level baselines with access to privileged information like expert demonstrations.
(a) Maze
(b) Pick and place
(c) Bin
(d) Kitchen
Figure 3: Non-stationarity metric comparison This figure compares the average distance metric between the subgoals predicted by the higher level policy and the states achieved by the lower level policy during training. As can be seen, LGR2 consistently generates efficient and achievable subgoals, thereby mitigating non-stationarity in HRL.
(a) Maze navigation
(b) Pick and place
(c) Bin
(d) Kitchen
@@
(a) Pick Start
(b) Pick Final
(c) Bin Start
(d) Bin Final
Figure 5: Real world experiments in pick and place and bin environments. Columns 1 and 2 depict initial and final row configurations for pick task respectively, and Columns 3 and 4 depict initial and final row configurations for bin task respectively.
Ablation Analysis We now analyze whether LGR2 indeed mitgates non-stationarity in HRL. In Figure 3, we compare LGR2 with vanilla HIER baseline, by computing the average distance between the subgoals predicted by the higher-level policy and the states achieved by the lower-level policy during training. A lower distance value implies that the higher level policy generates achievable subgoals, which induces optimal goal-reaching behaviour by the lower level policy, thereby mitigating non-stationary in HRL. Our findings reveal that LGR2 consistently maintains low average distances, thereby reducing non-stationarity. As seen in Figure 3, LGR2 achieves significantly lower distance values than the HIER baseline, demonstrating that it generates feasible subgoals.
Further, we analyse the importance of hindsight experience replay (HER) (Andrychowicz et al., 2017) for densifying rewards in sparse reward tasks. To this end, we compare LGR2 with LGR2-No-HER baseline, which is LGR2 baseline without HER. As seen in Figure 4, HER is crucial for good performance, since language-guided rewards are too sparse to generate any meaningful reward signal. We provide qualitative task visualizations in Supplementary Section F.
Real World Experiments We also perform real world experiments on pick and place and bin and environments (Fig 5). We use the Realsense D435 depth camera to extract the robotic arm, block and bin positions. We firstly train the hierarchical policies in simulation, and then deploy on real robotic arm. Computing accurate linear and angular velocities is hard in real tasks, so we assign them small hard-coded values, which shows good performance. We performed $5$ sets of experiments with $10$ trial each, and directly report the average success rates. LGR2 achieves an accuracy of $0.6$ and $0.5$, with variances $0.07$, $0.03$ on pick and place and bin tasks respectively. We also deployed the next best performing baseline L2R on the two tasks, but it failed to show any significant progress. In the supplementary, we also provide a detailed video depicting qualitative real-world results.
@@ 6 Conclusion
Limitations and future work Although we demonstrate that LGR2 is able to solve complex robotic control tasks through natural language instructions, it faces certain limitations. While we have tried to design complex sparse robotic control task environments which require complex reasoning, we do plan to explore more complex, multi-step tasks which might prove to be challenging for language based control in the future. Further, we use natural language instructions to generate stationary rewards for the higher-level policy, thus mitigating HRL non-stationarity. However, designing subgoals that are hard to express in language, like teaching a robot to do a cartwheel, remains challenging. A possible approach is to generate such complex subgoals directly from language while keeping higher-level rewards stable, which we plan to explore further in future work. Also, designing prompts for the reward translator module requires additional manual effort. Alternatively, if we neglect to design the prompts, the LLMs might hallucinate the parameters for the reward function. An interesting direction for future work is to automatically generate the prompts or even disregard the prompts altogether. Finally, in this work, we focus on solving the non-stationarity due to non-stationary higher rewards. However, we also plan to deal with non-stationary higher-level state transition functions in the future.
@@
Supplementary Materials
Appendix A Broader Impact Statement
Our proposed approach and algorithm are not aimed at delivering immediate technological breakthroughs. Instead, our work makes a conceptual contribution by addressing fundamental aspects of Hierarchical Reinforcement Learning (HRL). By leveraging language-guided instructions, we establish a novel framework with the potential to significantly propel research in HRL and related areas. This conceptual foundation lays the groundwork for future studies and may catalyze further advancements in the field.
Appendix B Implementation details
We conduct our experiments on two systems, each equipped with an Intel Core i7 processor, 48GB of RAM, and an Nvidia GeForce GTX 1080 GPU. We also report the number of timesteps required for running the experiments.
In our setup, both the actor and critic networks are implemented as three-layer, fully connected neural networks, each with 512 neurons per layer.
For the maze navigation task, a 7-degree-of-freedom (7-DoF) robotic arm moves through a four-room maze with its closed gripper fixed at table height, navigating to reach the goal position. In the pick-and-place task, the same 7-DoF robotic arm identifies a square block, picks it up, and delivers it to the goal position. In the bin environment, the gripper must pick up the block and place it in a designated bin. Lastly, in the kitchen task, a 9-DoF Franka robot performs a predefined complex action—opening a microwave door—to complete the task.
To ensure fair comparisons, we maintain consistency across all baselines by keeping key parameters unchanged wherever possible. These include the neural network layer width, the number of layers, the choice of optimizer, and the SAC implementation parameters.